### v1

# DocNet: Semantic Structure in Inductive Bias Detection Models

Abstract
News will have biases so long as people have opinions. However, as social media becomes the primary entry point for news and partisan gaps increase, it is increasingly important for informed citizens to be able to identify bias. People will be able to take action to avoid polarizing echo chambers if they know how the news they are consuming is biased. In this paper, we explore an often overlooked aspect of bias detection in documents: the semantic structure of news articles. We present DocNet, a novel, inductive, and low-resource document embedding and bias detection model that outperforms large language models. We also demonstrate that the semantic structure of news articles from opposing partisan sides, as represented in document-level graph embeddings, have significant similarities. These results can be used to advance bias detection in low-resource environments. Our code and data are made available at https://github.com/nlpresearchanon.
1 Introduction
The news media ecosystem has transformed dramatically in the past decade. Where previously, only established, professional news sources could easily disseminate information, now any outlet with an internet connection and social media presence can share their information to a broad audience Rodrigo-Ginés et al. (2023). Social media has further allowed not just the proliferation of fake news, but also the exacerbation of polarized echo chambers, where users are primarily exposed to and interact with media that already aligns with their biases Cinelli et al. (2021). As such, to maintain a well-informed society, where readers can be warned of biased reporting Spinde et al. (2022a) and can seek non-partisan or diverse reporting, it is imperative for resources for bias detection to be available.
However, given the inherently subjective nature of bias and the many different forms in which biases manifest in the media Rodrigo-Ginés et al. (2023), bias detection remains an inexact practice. Various datasets have been developed in the past five years to allow for greater research of biases at the word, sentence, article, and document level Zou et al. (2023); Wessel et al. (2023); Fan et al. (2019); Vallejo et al. (2023); Spinde et al. (2021a); Kiesel et al. (2019); Carragher et al. (2024). In addition, while automatic bias detection models historically relied on curated linguistic feature-based characterizations of text Kiesel et al. (2019); Spinde et al. (2021b), more recent studies have sought to use generalizable embeddings through deep learning-based methods, like contrastive learning Michelle YoungJin Kim (2022), pre-trained language models (e.g. BERT) Krieger et al. (2022); Vargas et al. (2023); Spinde et al. (2022b); Lei et al. (2022); Baly et al. (2020); Raza et al. (2022), LSTMs (Long Short-Term Memory neural networks) Roy and Goldwasser (2020); Lei et al. (2022), and Large Language Models (LLMs) Lin et al. (2024).
Although these deep learning methods are able to leverage word context, they do not factor in variances in writing style structures in media. As such, we propose an inductive, graph-based approach to partisan bias detection. We construct word co-occurrence graphs for each news article and find that their graph embeddings have comparable capture of bias as language model embeddings. Our detection model also statistically outperforms LLM-generated bias predictions. In this paper, our contributions are threefold:
  1. 1.
    We provide a methodology, DocNet, for a novel, low resource, and inductive document network and bias detection model that is not reliant on pre-trained language models.
  2. 2.
    We demonstrate the value of semantic structure in bias detection within the media ecosystem.
  3. 3.
    We identify high semantic structural similarity between news on either side of the political spectrum.
2 Background
Bias detection is a difficult problem, not least of all because of the many different forms of bias. Recent work by Rodrigo-Ginés, et al. has identified at least 17 different categories of bias Rodrigo-Ginés et al. (2023). Bias detection in the media is complicated due to the high level of text re-use in published news articles Boumans et al. (2018). A large part of widely circulated online news consists of redistributed or edited copies from large news agencies. This text re-use and the decisions behind what to produce or what to edit often result in instances of informational and lexical biases. Informational biases occur when news articles shape how information is perceived based on what surrounding or background information is shared in the article, and lexical bias occurs when polarizing words or clauses are used Fan et al. (2019). Identification of informational bias in a high text re-use media ecosystem needs to be able to isolate these potentially minute shifts in the structure of an article from otherwise similar but differently biased text.
News Bias Detection
Current news bias detection models use pre-trained word embeddings, like GloVe or Word2Vec, and language models, like BERT, to capture shifts in language in news articles. For instance, Zou et al., Zou et al. (2023) construct an article bias detection model using RoBerta-base and the POLITICS language model, which is a finetuned language model for ideology prediction Liu et al. (2022). Truica et al. combine TF-IDF (term frequency-inverse document frequency), and both non-pre-trained and pre-trained language model word embeddings to represent documents for article bias detection Truică and Apostol (2023). Meanwhile, Chen et al., use the BASIL dataset Fan et al. (2019) and predict article level biased based on the sentence level biases’ frequency, position, and sequential order Chen et al. (2020). They find that for the small BASIL dataset, lexical information (e.g. through bag of words), was insufficient to detecting article-level bias, while the features that characterize the internal structure of an article were more effective. Unfortunately, these methods require either a corpus of meticulously extracted sentence level and event features or a pre-trained language model.
Document Classification
Article-level bias detection can also be reframed as a document classification problem. Within the more generalized document classification space, methods remain similar, with most models highly dependent on pre-trained language models or expert generated features. Doc2vec Le and Mikolov (2014) is one of the few methods that requires neither. However, transformer based approaches where word embeddings from pre-trained language models are aggregated have largely surpassed Doc2vec’s performance. Document classification researchers have begun exploring aggregating these embeddings via graph representations Wang et al. (2023); Liu et al. (2021). Much of graph based classification research has been transductive, where one graph is generated for the entire corpus of documents, connecting word nodes across documents and generating document nodes for classification Zhao and Song (2023); Dai et al. (2022); Nikolentzos et al. (2020); Yao et al. (2018); Carragher et al. (2024). Transductive models, however, are not representative of reality and are not as computationally efficient as inductive models since they require one large graph to be retrained with every addition to the corpus. InducT-GCN is an inductive classification model that creates one graph per document using document node vectors and word node vectors. Word-word edges are connected by their pointwise mutual information (PMI), word-doc edges are connected using TF-IDF, and word embeddings use pre-trained embeddings Wang et al. (2022). They train a graph convolutional network (GCN) model for text classification. However, inductive document level models struggle with longer documents due to the dense graphs that are created Piao et al. (2022). All of these models also use graphs as a method to aggregate word embeddings in a document. They do not isolate semantic structure in the constructed graphs themselves.
3 Method
In this section, we detail our methodology for detecting the biases of online news articles. The training pipeline for our supervised bias detection models has four procedures, 1) Dataset collection, 2) Graph construction, 3) Unsupervised embedding generation and 4) Supervised model training.
3.1 Datasets
Our analysis uses four datasets, three of which are centered around partisan topics: withdrawal from Afghanistan, the Oath Keepers and the military vaccine mandate, and the fourth of which is the BASIL benchmark dataset Fan et al. (2019). (For brevity, we will refer to them as the AFG, OATH, VAX, and BASIL datasets, respectively moving forward). Table 1 summarizes the four datasets.
Table 1: Summary of datasets (Notes: “Far-Left” labels do not exist in this data. “Sentences” is shortened to “sent”)
Dataset AFG OATH VAX BASIL
# of Articles 5490 1965 1367 300
# of Domains 746 417 400 3
Sent/Article 39.3 40.4 37.0 26.6
Words/Sent 23.0 24.1 23.4 28.1
%Left 8.3 20.5 3.9 28.0
%Left-Center 40.8 48.0 23.6 00.7
%Center 20.2 23.8 26.1 46.0
%Right-Center 14.9 4.4 12.3 0.0
%Right 15.4 3.2 26.1 25.3
% Far Right 0.3 0.2 3.7 0.0
Topic Aligned Datasets
We introduce three, new, publicly available topic-based datasets. They are collected by first pulling Tweets from 1 February 2022 until 29 December 2022 on relevant news stories based on a keyword search via the Twitter API. Table 2 displays the collection terms.
Table 2: Collection Terms for Each Dataset
Collection Terms Dataset
{afghanistan, taliban, kabul, bagram AFG
U.S. withdrawal, afghan}
{oath keepers} OATH
{vaccine, vax} VAX
We extracted the URLs from Tweets and filtered out the social media site URLs, as has been done in other works Cruickshank and Carley (2020, 2021). From there, we scraped the textual content of those websites that allowed automated scraping. This methodology results in a variety of news sources and articles that are relevant to discussions on these partisan topics Cruickshank et al. (2023). Weak article based bias labels (far left, left, left-center, center, right-center, right, far right) for each article are assigned based upon its domain bias label in lieu of human annotations Cruickshank and Carley (2021). This follows the assumption that for highly partisan topics, the level of bias of a news article and the events they discuss will be highly correlated with the partisan sentiment of its domain Zou et al. (2023). We further validate the reliabiltiy of our domain labels by comparing them with the bias levels from Adfontes, as of 10 March 2023  [^1] https://adfontesmedia.com/. Per a Pearson correlation test of domain bias labels using an ordinal encoding, our labels are highly correlated with Adfontes’ with a correlation coefficient of .79 and p-value $<e^{-80}$.
BASIL
We further validate our approach with a human annotated article level dataset, the BASIL dataset. The BASIL dataset is composed of 100 articles on the same event from three different news outlets. Each article is manually annotated with article bias labels Fan et al. (2019).
3.2 Graph Construction
We transform each news article into its own graph in order to create article embeddings after the text is cleaned and lemmatized. Stop words are also removed. Each graph is an undirected word (or lemma) co-occurrence network. With these word nodes, an edge is drawn between co-occurring words, where the weight of the edge is the positive pointwise mutual information (PPMI), Equation 1. The edge is removed if there is a non-positive PMI.
$\begin{split}PMI=\frac{log_{2}(P_{w_{1}w_{2}}/(P_{w_{1}}*P_{w_{2}}))}{-log_{2}(P_{w_{1}w_{2}})})\\ P_{x}=\frac{\text{Frequency of $x$ in document}}{\text{Total Unigrams or Bigrams}}\\ \text{$w_{1}$, $w_{2}$ are the words that construct edge $(w_{1},w_{2})$}\end{split}$ (1)
We construct three variations of this graph
  1. 1.
    Base graph: No node attributes are provided
  2. 2.
    DocNet: The node attribute is the originating word, $w_{i}$
  3. 3.
    Node vector graph: The node attribute is the word vector per Spacy’s en_core_web_lg model  [^2] Version 3.4.1 Honnibal and Montani (2017) concatenated with the average sentiment of the given lemma across all occurrences in the article per NLTK VADER’s Sentiment Intensity Analyzer Hutto and Gilbert (2014).
3.3 Unsupervised Embeddings
Next, we generate unsupervised embeddings for each article. The embedding models are inductively trained using an 80%/20% train/test split. We do not conduct rigorous hyperparameter tuning, but rather seek to align parameters across models for better comparison.
Graph2Vec
We train three Graph2Vec models for each graph variation using the KarateClub package Rozemberczki et al. (2020). Graph2Vec conducts Weisfeiler-Lehman (WL) hashing to generate graph level features that are then passed as the “document” in Doc2Vec Le and Mikolov (2014). In the base graph’s case, the hashing step is conducted using node degrees, whereas the designated node attribute is used otherwise (e.g. the word, or the word vector from a pre-trained language model). The PPMI scores are not factored into the model as Graph2Vec does not support edge attributes. We set a dimension size of 128, 50 epochs, and 2 Weisfeiler-Lehman iterations.
Graph Auto-Encoders
We also train a graph auto-encoder model (“gae”) and a variational graph auto-encoder model (“vgae”) Kipf and Welling (2016) on the word vector graph representation of the text. Both graph auto-encoders are constructed using three modules of graph convolution, graph normalization, and rectified linear unit (ReLU) layers in the encoder and an inner product decoder. In order to have a comparable embedding size and epochs to the other models, we set output feature size to be 128. We also use an Adam optimizer with a starting learning rate of .1 and a scheduler that will reduce the learning rate when the loss plateaus after a minimum of five epochs and maximum of fifty epochs.
Doc2Vec
As a comparison to the graph based approaches, we also use Doc2Vec Le and Mikolov (2014), using both the distributed memory (“doc2vec_dm”) and the distributed bag of words (“doc2vec_dbow”) representations. We set a window size of 10, initial learning rate of .05, 50 epochs, minimum word count to be 3, negative sample of 5, and a vector size of 128
Doc2Vec and Graph2Vec
In order to measure if graph based embeddings capture anything different from Doc2Vec, we also concatenate the separately trained Graph2Vec and Doc2Vec embeddings for a 256 dimension sized vector in the final supervised model. We combine the two variants of Doc2Vec with the three variants of Graph2Vec for a total of six different embedding models.
Sentence BERT
As a final unsupervised embedding baseline comparison, we use Sentence BERT Reimers and Gurevych (2019). For each sentence in an article, we pass it through the pre-trained “All-Distilroberta-V1” encoder to map it to the default 768 dimensional vector space. A given article is subsequently represented by the average of all of its sentence embeddings.
3.4 Bias Detection
We explore the value of these various embeddings in partisan bias detection using a naive multinomial logistic regression model. We train three different inductive models using different bias label groupings 1) “Full”: the original bias categories, 2) “LCR”: Re-grouping mildly or extremely partisan articles into right or left, for three nominally encoded bias categories (right, center, left) and 3) “Binary”: Re-grouping into biased or unbiased, where left/right-center are both considered to be unbiased. After model predictions are generated, we further manually regroup the “Full” category labels into the “LCR” and “Binary” to more accurately compare accuracy between models with different numbers of categories. This allows us to explore the level of granularity in bias information in the various embeddings as well as the potential for similarities between opposing sides in partisan discussions.
Domain Aggregation
Due to the weak article bias labels used in the AFG, OATH, and VAX datasets, we also explore training domain level bias detection models. There is a varied number of articles between domains in the topic aligned dataset, so we aggregate the embeddings for a given domain (i.e. news source), and train a bias detector based on the combined representation of the de-duplicated domains’ article embeddings. There is only one sample of each domain after articles are aggregated. We aggregate articles in a given domain by taking the arithmetic mean of the embeddings. We also explore the effect of normalizing the embeddings by its original topic. We subtract the average embeddings within a given topic from the embedding of each article and take the average (“avgTopicDiff”) or Frobenius norm of the difference between those articles (“froNormTopicDiff”). These topic based aggregation techniques aim to remove any topical skew in the data when all three topic based datasets are trained together. With the BASIL dataset, our domain level models predict the bias level of the domain for each article without aggregating across domains since there are only three different news sources.
Large Language Models
As a final comparison, we conduct document level bias detection using pre-trained large language models finetuned for chat completion, OpenAI’s GPT Turbo 3.5 OpenAI (2021) and Meta’s Llama 3 8B Instruct Model AI@Meta (2024). We use the following zero-shot prompt template to generate bias predictions: “Please classify the following news article by its political bias. Please only classify the article as {bias labels}, and return no other text. title: {title} article: {article} bias: ”. We do not conduct domain level predictions using the large language models.
Naive Baseline
We use a Zero Rule baseline for a naive model comparison where the predictions are all labeled as the majority class.
4 Results and Discussion
We ran 1284 experiments to train multiple embeddings and supervised learning models over the various configurations and four datasets, to include a combined topic agnostic dataset that merges the AFG, OATH, and VAX datasets. The experiments are trained on 1 node of an NVIDIA DGX A100 system. Each graph autoencoder embedding model takes at most 2 GPU hours to train one model, while each Graph2Vec embedding model takes at most 3.5 CPU minutes. In order to control for different configurations, we formally evaluate the difference between configurations and embeddings using a one-sided Wilcoxan signed rank test. This allows us to determine the paired difference between two methods when all other configurations are held constant. We are able to confirm that the embeddings are valuable in comparison to the naive baseline for all metrics, with p-values close to 0. Our experiments have an average of .21 increase in macro F-1 scores over the baseline. This is on average a 120% improvement over the baseline.
Table 3: Average (std) of results across all embedding types
Accuracy Macro F-1
Model
graph2vec_no_features 0.498528 (.169) 0.341580 (.137)
GPT 3.5 Turbo 0.416704 (.164) 0.349551 (.177)
VGAEautoencode_data_sent 0.474973 (.143) 0.357707 (.128)
VGAEautoencode_data 0.470229 (.155) 0.362188 (.135)
Llama 3-8B 0.453731 (.225) 0.378090 (.204)
GAEautoencode_data_sent 0.524193 (.146) 0.391000 (.135)
GAEautoencode_data 0.537793 (.146) 0.425286 (.138)
doc2vec_dbow_graph2vec_no_features 0.559085 (.169) 0.456914 (.170)
doc2vec_dm_graph2vec_no_features 0.543820 (.178) 0.466900 (.184)
doc2vec_dm 0.538900 (.170) 0.472554 (.170)
doc2vec_dbow 0.573575 (.170) 0.485918 (.170)
doc2vec_dm_graph2vec_nodewords 0.583128 (.164) 0.500872 (.167)
graph2vec_spacy 0.588646 (.154) 0.501290 (.161)
doc2vec_dbow_graph2vec_spacy 0.584941 (.163) 0.501290 (.165)
graph2vec_nodewords 0.591630 (.151) 0.506795 (.157)
doc2vec_dm_graph2vec_spacy 0.572459 (.157) 0.508879 (.160)
doc2vec_dm_graph2vec_nodewords 0.570328 (.164) 0.510572 (.167)
SBERT 0.606528 (.154) 0.541717 (.162)
Label Aggregation
The results of models grouped on different label bins are in Table 4. The large variation in scores is primarily a result of the increased difficulty with the greater number of classes. We find that after regrouping the “Full” label models’ predictions, there is not a statistically significant difference between models trained on the “Full” label set in comparison to “LCR” or “Binary” labels, p-value of .91 and .36, respectively. This demonstrates that while the “Full” multiclass models have inherently lower scores, they are able to extract divisions in the news’ embedding space just as successfully as the models trained on fewer categories.
Table 4: Average (std) of results based on different label aggregations
Accuracy Macro F-1 Baseline Accuracy Baseline Macro F-1
Label Aggregation
Full 0.403810 (.102) 0.293628 (.111) .379535 (.075) .105336 (.038)
LCR 0.517936 (.102) 0.465078 (.102) .470023 (.067) .216987 (.036)
Binary 0.722017 (.099) 0.608042 (.118) .7166878 (.069) .416615 (.024)
Domain Aggregation
The results for the different domain aggregation comparisons are listed in Table 5. We find that with the BASIL dataset, the embeddings separated more easily into domain level labels (although the article models were still significantly better than the naive baseline across all metrics with a p-value near 0). We also find that the method of aggregating article embeddings makes a significant difference on performance, such that domain models using the Frobenius norm of the difference from its topic perform worse than news level predictions. The mean of article embeddings across all samples of a domain has the best performance in domain level bias detection.
Table 5: Average (std) of results across additional configurations
Accuracy Macro F-1
Domain Aggregation
froNormTopicDiff 0.500561 (.159) 0.397489 (.153)
no_aggregation 0.564704 (.165) 0.453129 (.167)
avgTopicDiff 0.555718 (.163) 0.483041 (.167)
arithmeticMean 0.563010 (.168) 0.489796 (.175)
BASIL_articles 0.372284 (.111) 0.332773 (.112)
BASIL_domains 0.587901 (.190) 0.543800 (.196)
Article Subsets
no_filter 0.520002 (.173) 0.445451 (.174))
filterAfg 0.549137 (.157) 0.447124 (.143)
filterVaccine 0.510295 (.148) 0.450861 (.166)
filterMultArticles 0.535061 (.168) 0.452376 (.168)
filterMultTopic 0.578134 (.171) 0.463696 (.195)
4.1 Embedding Model
The average mean and standard deviation of scores given different embeddings and models are in Table 3. Through the paired rank test on the full set of experiments, we find that SBERT outperforms all other embeddings using a p-value significance threshold of .05. The GPT, Llama, graph autoencoders, and Graph2Vec with no features (i.e. node attributes as degrees) perform the worst, with no statistically significant performance difference between them. The unsupervised Graph2Vec based embeddings are able to capture characteristics of biased text with performance often times similar to, and in the case of “graph2vec_nodewords”, 30% of the time, better than SBERT. Graph2Vec with node attributes (both with just node words and word vectors) outperforms both doc2vec models with p-values all less than .01. We also observe that when Graph2Vec embeddings with node features are combined with doc2vec, a statistically significant improvement (p-values are all less than .01) from the isolated doc2vec model occurs. For instance, Graph2Vec with node words concatenated with doc2vec with distributed memory (“doc2vec_dm_graph2vec_nodewords”) consistently outperforms embeddings using just “doc2vec_dm”, as observed in the their average scores in Table 3. There is no statistical difference between the type of node attribute used in Graph2Vec (node words or Spacy vectors).
These results demonstrate the value of including semantic structure in bias classification so long as there is a unique node identifier that can be used to indirectly relate separate graphs. Our Graph2Vec models do not factor in edge attributes, and its only information on word frequency is the lack of certain edges due to non positive PMI. Unlike with SBERT, language models and word embeddings also are not necessary in this characterization. An indirect tie between document networks is necessary and sufficient to have performance that often times is comparable to transformer based document embeddings. Therefore, characterizing text using individual co-occurrence graphs results in improved bias detection.
While our computationally expensive graph autoencoder models pass both edge and node vector attributes in generating document embeddings, they still under perform in comparison to the simple Graph2Vec unsupervised embeddings. Further experimentation and tuning should be conducted to validate if alternative deep learning graph models, like graph attention or diffusion based approaches Nikolentzos et al. (2020); Liu et al. (2021), could be repurposed for inductive embedding generation, as graph neural network based approaches are often prone to oversmoothing. Regardless, the fact that a simple approximate graph isomorphism comparison through WL-hashing with minimal additional attributes can generate good bias predictions, validates the value of characterizing semantic structure using document networks.
4.2 DocNet Characteristic Case Study
We finally explore if the models that use Graph2Vec embeddings are aligning along trivial graph features. In Figure 1 we plot graphs from the BASIL dataset using a spring force-directed algorithm [^3] https://networkx.org/documentation/stable/reference/generated/networkx.drawing.layout.spring_layout.html. We select articles that have the highest confidence (i.e. probability) predictions for the respective incorrect or correct binary bias labels from a Graph2Vec with node words embedding. We observe that in Figure 1, the graphs for the articles that were predicted to be unbiased (a and d) both have visually sparse layouts. Meanwhile, the biased articles (b and c) appear to be denser. Upon returning to the original data, we also find that the news article from Figure 1d is actually labeled at the article level to be unbiased. However, the original article label for Figure 1c is unbiased. This article  [^4] https://archive.nytimes.com/www.nytimes.com/politics/first-draft/2016/02/25/2-ex-presidents-of-mexico-say-no-way-theyre-.paying-for-donald-trumps-wall contains phrases like “Donald J. Trump leads a familiar refrain” and “Mr. Trump, who has been criticized for lacing his campaign speeches with profanity”. This causes us to question the validity of this article’s human annotated bias label. Furthermore, these examples demonstrate the continued difficulty in developing good labels for an inherently subjective application. Yet these graph plots do imply that there is a pattern in the graph structure that captures bias.
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-Pred Biased
(d) Incorrect-Pred Unbiased
Figure 1: Graphs of articles with high probability predictions from BASIL. Predictions from graph2vec_nodewords at the domain level and binary labels (Macro F-1 = .82). Source domains: a) NYTimes, b) Fox, c) NYTimes d) Huffington Post
Figure 2: Correlation Plot of Graph Metrics and Bias from Topic Aligned Dataset
Figure 3: Correlation Plot of Graph Metrics and Bias from BASIL (Art is article bias, D is domain bias)
To further explore the characteristics of the semantic structure, we obtain various descriptive graph metrics for each article and calculate their correlation with bias labels as shown in Figures 2 and 3. In both correlation plots, there are no strong correlations between any of the metrics and the corresponding bias labels. We then train a simple logistic regression model on these statistics for each dataset to predict binary bias labels. These models have a .21 lower Macro F-1 on the partisan dataset and a .12 lower Macro F-1 on the domain level Basil predictions compared to the corresponding Graph2Vec with node words model. This analysis further demonstrates that while the semantic structure holds bias information, there is additional information to be gained by using our proposed graph embeddings rather than curated descriptive graph metrics.
5 Conclusion and Future Work
Through our graph embeddings we demonstrate that encoding semantic structure in news articles as co-occurrence networks can be used for both domain and article level bias detection. We couple baseline embeddings with a logistic regression to learn a separation of the embedding space and learn that bias texts have similar word context agnostic semantic structures that can be captured by document networks (i.e. DocNet). This is demonstrated by the success of the Graph2Vec embeddings in the supervised models, especially as it often has comparable performance to the SBERT embeddings, even without the use of word embeddings. DocNet does not require pre-trained language models to generate word embeddings, nor does it require computationally expensive deep learning techniques. The most computationally time consuming portion of DocNet is the initial graph construction of an article. However, this can be easily parallelized across a corpus of documents. It is completely inductive, allowing for inexpensive prediction on new datasets.
Through the various groupings of bias labels, we also identify that there is often insufficient information in the semantic structure of an article for accurate fine-grained bias classification, although future research using one-vs-rest classifiers rather than our generalized multinomial regression may be able to improve upon our results. This is an observation observed by other recent work Carragher et al. (2024). On the other hand, fuzzier binary classifications can still be sufficiently characterized by semantic structures. This means that right-leaning news articles and left-leaning news articles have significantly more in common than either side of the political spectrum may believe. Although we are not able to pinpoint exactly what this writing style bias is due to the lack of interpretability in our graph embeddings, we can definitively conclude that the characteristics of partisan rhetoric are more than context-dependent.
In future work, it would be valuable to research embeddings that can fuse word embeddings with semantic structure. While our graph autoencoders performed better than a naive baseline, they were significantly less useful than our Graph2Vec embeddings in bias detection. Context represented in pre-trained language models is valuable to bias detection, as demonstrated in our SBERT model. Future developments in inductive graph embedding approaches that can efficiently and effectively represent dense networks and attributes would improve bias detection efforts. In the interim, our DocNet methodology provides a pathway for low-resourced bias detection, where computational resources, rigorously labeled datasets, or language models are not available.
6 Limitations
This project primarily relies on partisan topics and does not look into bias detection along non-partisan spectrums. In addition, as the data was collected based on what Tweets on the same topic were shared, there may be a skew in our dataset to highlight media with characteristics that draw the attention of Twitter users. However, given the size of the dataset, the variation in domains present, and the consistency in trends with the manually labeled BASIL dataset, we believe our data to be sufficiently generalizable. Without further data validation, our data should only be used for research purposes. An incorrect bias detection model may lead to misplaced mistrust or over confidence in media. This project also only analyzes articles in the English language. While we believe our methodology is language agnostic, this hypothesis has not yet been validated on news articles from other languages.
Disclaimer
The views expressed in this paper are those of the authors and do not reflect the official policy or position of the United States Military Academy, the United States Army, the Department of Defense, or the United States Government.
7 Ethics Statement
All articles collected for this study were done so under the provisions of Section 107 of the U.S. Copyright Act and ensured that our collection action fell under the fair use category. The Tweets were collected in accordance with Twitter’s terms of service at the time of collection. The code and news articles are made available at https://github.com/nlpresearchanon/DocNet under the MIT and CC0 1.0 Universal licenses, respectively.