### v2

# DocNet: Semantic Structure in Inductive Bias Detection Models

Abstract
News will have biases so long as people have opinions. It is increasingly important for informed citizens to be able to identify bias as social media becomes the primary entry point for news and partisan differences increase. If people know the biases of the news they are consuming, they will be able to take action to avoid polarizing echo chambers. In this paper, we explore an often overlooked aspect of bias detection in documents: the semantic structure of news articles. We present DocNet, a novel, inductive, and low-resource document embedding and bias detection model that outperforms large language models. We also demonstrate that the semantic structure of news articles from opposing partisan sides, as represented in document-level graph embeddings, have significant similarities. These results can be used to advance bias detection in low-resource environments. Our code, data, and the corresponding datasheet are made available at:
Code and Datahttps://anonymous.4open.science/r/DocNet/
1 Introduction
The news media ecosystem has transformed dramatically in the past decade. Where previously only established, professional news sources could easily disseminate information, now any outlet with an internet connection and social media presence can share their information to a broad audience (Rodrigo-Ginés, de Albornoz, and Plaza 2023). Social media has allowed not just the proliferation of fake news, but also the exacerbation of polarizing echo chambers, where users are only exposed to and interact with media that already aligns with their biases (Cinelli et al. 2021). As such, to maintain a well-informed society, where readers can be warned of biased reporting (Spinde et al. 2022a) and can seek non-partisan or diverse reporting, it is imperative that resources for automatic article level bias detection to be available.
Given the inherently subjective nature of bias and the many different forms in which biases manifest in the media (Rodrigo-Ginés, de Albornoz, and Plaza 2023), bias detection remains an inexact practice. Various datasets have been developed in the past five years to allow for increased research of biases at the word, sentence, article, and document level (Zou et al. 2023; Wessel et al. 2023; Fan et al. 2019; Vallejo, Baldwin, and Frermann 2023; Spinde et al. 2021a; Kiesel et al. 2019; Carragher, Williams, and Carley 2024). These automatic bias detection models have historically relied on curated linguistic feature-based characterizations of text (Kiesel et al. 2019; Spinde et al. 2021b). However, more recent studies have sought to use generalizable embeddings through deep learning-based methods, like contrastive learning (Kim and Johnson 2022), pre-trained language models (Krieger et al. 2022; Vargas et al. 2023; Spinde et al. 2022b; Lei et al. 2022; Baly et al. 2020; Raza, Reji, and Ding 2022), LSTMs (Long Short-Term Memory neural networks) (Roy and Goldwasser 2020; Lei et al. 2022), and Large Language Models (LLMs) (Lin et al. 2024).
Although these deep learning methods leverage word context, they do not explicitly factor in variances in writing style structures in media. As such, we propose an inductive, graph-based approach to partisan bias detection. We construct word co-occurrence graphs for each news article and discover that their graph embeddings have comparable performance in capturing partisan bias as language model embeddings. Our detection model also statistically outperforms LLM-generated bias predictions. In this paper, our contributions are threefold:
  1. 1.
    We provide a methodology, DocNet, for a novel, low resource, and inductive document network and bias detection model that is not reliant on pre-trained language models.
  2. 2.
    We demonstrate the value of semantic structure in bias detection within the media ecosystem.
  3. 3.
    We identify high semantic structural similarity between news on either side of the political spectrum.
2 Background
Bias detection is a difficult problem, not least of all because of the many different forms of bias. Recent work by Rodrigo-Ginés, et al. has identified at least 17 different categories of bias (Rodrigo-Ginés, de Albornoz, and Plaza 2023). Bias detection in the media is further complicated due to the high level of text re-use in published news articles (Boumans et al. 2018) and the different frames available to discuss an issue (Card et al. 2015). A large part of widely circulated online news consists of redistributed or edited copies from large news agencies. This text re-use and the decisions behind what to publish or how to edit an article result in instances of informational and lexical biases. Informational biases occur when news articles change the surrounding or background information in order to shape how information is framed, and thereby perceived. Lexical bias occurs when polarizing words or clauses are used to illicit specific emotion towards an issue (Fan et al. 2019). Partisan biases in news articles are generated through these shifts in content and languages. Identification of these rhetorical methods in a high text re-use media ecosystem requires isolating minute shifts in the language and structure of otherwise similar articles.
News Bias Detection
Current news bias detection models use pre-trained word embeddings, like GloVe or Word2Vec, and language models, like BERT, to capture biased language in news articles. For instance, Zou et al., (Zou et al. 2023) construct an article bias detection model using RoBerta-base and the POLITICS language model, which is a finetuned language model for ideology prediction (Liu et al. 2022). Truica et al. combine TF-IDF (term frequency-inverse document frequency), and both non-pre-trained and pre-trained language model word embeddings to represent documents for article bias detection (Truică and Apostol 2023). Meanwhile, Chen et al., use the BASIL dataset (Fan et al. 2019) and predict article level biases using the sentence level biases’ frequency, position, and sequential order (Chen et al. 2020). They find that for the small BASIL dataset, lexical information (e.g. through bag of words), was insufficient in detecting article-level bias, while the features that characterize the internal structure of an article were more effective. Yet, all of these methods hinge on either a corpus of meticulously extracted sentence level and event features or a suitable pre-trained language model.
Document Classification
Article-level bias detection can also be reframed as a document classification problem. Within the more generalized document classification space, methods remain similar, with most models highly dependent on pre-trained language models, like Longformers (Beltagy, Peters, and Cohan 2020) or expert generated features. Doc2vec (Le and Mikolov 2014) is one of the few methods that requires neither. However, transformer based approaches, where word embeddings from pre-trained language models are aggregated, have largely surpassed Doc2vec’s performance.
Document classification researchers have also begun aggregating these embeddings via graph representations (Wang, Ding, and Han 2023; Liu et al. 2021). Much of graph based document classification research has been transductive, where one graph is generated for the entire corpus of documents. Where word nodes are connected across documents and document nodes are generated for classification (Zhao and Song 2023; Dai et al. 2022; Nikolentzos, Tixier, and Vazirgiannis 2020; Yao, Mao, and Luo 2019; Carragher, Williams, and Carley 2024). Transductive models, however, are not representative of reality and are not as computationally efficient as inductive models since they require one large graph to be retrained with every addition to the corpus.
InducT-GCN is one of the few inductive document classification models discussed in literature (Wang, Han, and Poon 2022). It creates one graph for the entire training corpus using document node vectors and word node vectors. Word-word edges are connected by their pointwise mutual information (PMI) and word-doc edges are connected using TF-IDF. They train a graph convolutional network (GCN) model to learn the word embeddings and conduct classification. These word embeddings and learned weights are used to generate predictions on batches of test documents. However, other researches have found that inductive document level models struggle with longer documents due to the dense graphs that are created (Piao et al. 2022). This is apparent in the case of Induct-GCN, where while still outperforming baselines, there is observably worse performance on longer text document datasets like 20 News Group and Ohsumed, than for shorter ones like R52 (Wang, Han, and Poon 2022). The aforementioned graph based models have also only used graphs as a method to aggregate word embeddings in a document. They do not isolate the semantic structure of individual documents in the constructed graphs.
3 Method
In the following section, we detail our methodology for detecting the biases of online news articles. The pipeline for our inductive supervised bias detection models has four procedures, 1) Dataset collection, 2) Graph construction, 3) Unsupervised embedding generation and 4) Supervised model training.
Datasets
Our analysis uses four datasets, three of which are centered around partisan topics: withdrawal from Afghanistan, the Oath Keepers and the military vaccine mandate, and the fourth of which is the BASIL benchmark dataset (Fan et al. 2019). (For brevity we will refer to them as the AFG, OATH, VAX, and BASIL datasets, respectively, moving forward).
Topic Aligned Datasets
We introduce three publicly available topic-based datasets, the AFG, Oath, and VAX news article datasets. We chose these three topics for experimentation due to the significant attention, controversy, and corresponding media coverage surrounding each. Each topic’s respective articles are collected by first pulling Tweets from 1 February 2022 until 29 December 2022 on relevant news stories based on a keyword search via the Twitter API. Table 1 displays the collection terms.
Collection Terms Dataset
{afghanistan, taliban, kabul, bagram AFG
U.S. withdrawal, afghan}
{oath keepers} OATH
{vaccine, vax} VAX
Table 1: Collection Terms for Each Dataset
We extracted the URLs from Tweets and filtered out the social media site URLs, as has been done in other works (Cruickshank and Carley 2020, 2021). From there, we scraped the textual content of those websites that allowed automated scraping. This methodology results in a variety of news sources and articles that are relevant to discussions on these partisan topics (Cruickshank, Zhu, and Bastian 2023). We apply domain bias labels from a dataset compiled by Cruickshank et al (Cruickshank and Carley 2021) and available under CC by 4.0 licensing. Domains are the publishing sites that news articles are scraped from. We use the term news domain in lieu of news sources, as articles are often shared from distributing news agencies, like the Associated Press, to multiple sites (i.e. domains).
A weak bias label (far left, left, left-center, center, right-center, right, far right) for each article is assigned based upon its domain’s bias label. Our usage of domain labels as weak labels for article bias follows the assumption that for highly partisan topics, the bias of a news article and the events that are discussed are highly correlated with the perspective of its publisher, or domain (Zou et al. 2023). We further validate the reliability of Curickshank et al’s domain bias labels by comparing them with the domain bias levels published on Adfontes, as of 10 March 2023  [^1] https://adfontesmedia.com/. Per a Pearson correlation test of news domain bias labels using an ordinal encoding, we validate that our labels are highly correlated with Adfontes’ with a correlation coefficient of .79 and p-value $<e^{-80}$. This is in spite of the three year difference in times at which these domain bias labels were collected (2020 vs 2023) and the likely drift in media bias over time. This test increases confidence that the domain bias labels we use have minimal bias themselves, as both data sources report similar labels.
BASIL
We further validate our approach with a dataset that has human annotated article level biases, the BASIL dataset (Fan et al. 2019). The BASIL dataset is composed of 300 articles from 2010 to 2019 that cover 100 events, with one article each from Fox News, New York Times, and Huffington Post. Each article is human annotated with article and domain bias labels. Key characteristics of these four datasets are summarized in table 2.
Dataset AFG OATH VAX BASIL
# of Articles 5490 1965 1367 300
# of Domains 746 417 400 3
Sent/Article 39.3 40.4 37.0 26.6
Words/Sent 23.0 24.1 23.4 28.1
%Left 8.3 20.5 3.9 28.0
%Left-Center 40.8 48.0 23.6 00.7
%Center 20.2 23.8 26.1 46.0
%Right-Center 14.9 4.4 12.3 0.0
%Right 15.4 3.2 26.1 25.3
% Far Right 0.3 0.2 3.7 0.0
Table 2: Summary of datasets (Note: “Far-Left” labels do not exist in this data. “Sentences” is shortened to “sent”)
Graph Construction
Next, we transform each news article into its own graph after the text is cleaned and lemmatized in order to capture each article’s individual semantic structure in the article’s embedding. Each graph is an undirected word (or lemma) co-occurrence network. With these word nodes, an edge is drawn between co-occurring words, where the weight of the edge is the positive pointwise mutual information (PPMI), Equation 1. The edge is removed if there is a non-positive PMI.
$\begin{split}PMI=\frac{log_{2}(P_{w_{1}w_{2}}/(P_{w_{1}}*P_{w_{2}}))}{-log_{2}(P_{w_{1}w_{2}})})\\ P_{x}=\frac{\text{Frequency of $x$ in document}}{\text{Total Unigrams or Bigrams}}\\ \text{$w_{1}$, $w_{2}$ are the words that construct edge $(w_{1},w_{2})$}\end{split}$ (1)
We reserve the cleaned text of the article for comparison model experimentation and construct three variations of the article graph .
  1. 1.
    Base graph: No node attributes are provided
  2. 2.
    DocNet: The node attribute is the originating word, $w_{i}$
  3. 3.
    Node vector graph: The node attribute is the word vector per Spacy’s en_core_web_lg model  [^2] Version 3.4.1 (Honnibal and Montani 2017) and concatenated with the average sentiment of the given lemma across all occurrences in the article per NLTK VADER’s Sentiment Intensity Analyzer (Hutto and Gilbert 2014).
Unsupervised Embeddings
From the preprocessed texts, we generate unsupervised embeddings of each article. The embedding models are inductively trained using an 80%/20% train/test split. We do not conduct hyperparameter tuning. Instead, we seek to generally align parameters, like epochs and dimension size, across models for better comparison, or otherwise use default package hyperparameters.
Graph2Vec
We train three Graph2Vec models for each graph variation using the KarateClub package (Rozemberczki, Kiss, and Sarkar 2020). Graph2Vec conducts Weisfeiler-Lehman (WL) hashing to generate graph level features that are then passed as the “document” in Doc2Vec (Le and Mikolov 2014). In the base graph’s case, the hashing step is conducted using node degrees, whereas the designated node attribute is used otherwise (e.g. the word, or the word vector from a pre-trained language model). The PPMI scores are not factored into the model as Graph2Vec does not currently support edge attributes. We set a dimension size of 128, 50 epochs, and 2 Weisfeiler-Lehman iterations.
Graph Auto-Encoders
We also train a graph auto-encoder model (“gae”) and a variational graph auto-encoder model (“vgae”) (Kipf and Welling 2016) on the node vector graph representation of the text, both with and without the sentiment vector (“sent”). Both graph auto-encoders are constructed using three modules of graph convolution, graph normalization, and rectified linear unit (ReLU) layers in the encoder, and an inner product decoder. In order to have a comparable embedding size and epochs to the other models, we set output feature size to be 128. We use an Adam optimizer with a starting learning rate of .1 and a scheduler that reduces the learning rate when the loss plateaus after a minimum of five epochs and maximum of fifty epochs.
Doc2Vec
As a comparison to the graph based approaches, we use Doc2Vec (Le and Mikolov 2014), using both the distributed memory (“doc2vec_dm”) and the distributed bag of words (“doc2vec_dbow”) representations. We set a window size of 10, initial learning rate of .05, 50 epochs, minimum word count to be 3, negative sample of 5, and a vector size of 128.
Doc2Vec and Graph2Vec
In order to measure if graph based embeddings capture anything different from Doc2Vec, we also concatenate the separately trained Graph2Vec and Doc2Vec embeddings into a 256 dimension sized vector in the final supervised model. We combine the two variants of Doc2Vec with the three variants of Graph2Vec for a total of six different embedding models.
Sentence BERT
As a final unsupervised embedding comparison, we use Sentence BERT (Reimers and Gurevych 2019). For each sentence in an article, we pass it through the pre-trained “All-Distilroberta-V1” encoder to map it to a default 768 dimensional vector space. A given article is subsequently represented by the average of all of its sentence embeddings.
Bias Detection
We explore the value of these various embeddings in partisan bias detection using a naive multinomial logistic regression model. We train three inductive models using varied bias label groupings 1) “Full”: the original bias categories, 2) “LCR”: Re-grouping mildly or extremely partisan articles into right or left, for three nominally encoded bias categories (right, center, left) and 3) “Binary”: Re-grouping into biased or unbiased, where left/right-center are both considered to be unbiased. After predictions are generated, we further manually regroup the “Full” category labels into the “LCR” and “Binary” to more accurately compare accuracy of models with different numbers of categories. These variations allow us to explore the level of granularity of the bias information captured in the various embeddings. It also allows us to interrogate the potential for similarities between opposing viewpoints in partisan discussions.
Domain Aggregation
We also explore training domain level bias detection models due to the weak article bias labels used in the AFG, OATH, and VAX datasets. In the topic aligned datasets, there is a mixed number of articles between domains. As a result, we develop various methods to deduplicate domains for training a news domain bias detector by aggregating the article embeddings of each domain. In other words, the domain bias detectors are trained on a combined representation of each domain’s article embeddings.
We apply three methods of aggregating articles for each domain. First, we aggregate articles in a given domain by taking the arithmetic mean of the embeddings (“arithmeticMean”, equation 2). We also explore the effect of normalizing the embeddings by its original topic. We subtract the average embeddings within a given topic from the embedding of each article and take the average (“avgTopicDiff”, equation 4), and for another method, take the Frobenius norm (“froNormTopicDiff”, equation 5). The latter two topic based aggregation techniques aim to remove any topic specific skew in the data when all three topic aligned datasets are aggregated.
arithmeticMean: $\begin{split}\text{arithmeticMean: }d_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}\sum_{t=1}^{n_{t}}x_{ijt}\\ \end{split}$ (2)
$\begin{split}T_{t}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}x_{ijt}\\ \hat{x}_{ijt}=x_{ijt}-T_{t}\end{split}$ (3)
avgTopicDiff: $\begin{split}\text{avgTopicDiff: }d_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}\sum_{t=1}^{n_{t}}\hat{x}_{ijt}\\ \end{split}$ (4)
froNormTopicDiff: $\begin{split}\text{froNormTopicDiff: }d_{j}=\sqrt{\sum_{i=1}^{n_{j}}|\hat{x}_{ijt}|^{2}}\\ \end{split}$ (5)
where,
$d_{j}=\text{embedding of domain, $j$}$
$x_{ijt}=\text{embedding of article $i$ from domain $j$ and topic$t$}$
$n=\text{number of articles}$
$T_{t}=\text{average topic embedding of topic $t$}$
With the BASIL dataset, our domain level models predict the bias level of the domain for each article without aggregating embeddings as there are only three different news domains. This increases the risk of the model capturing spurious characteristics, but in reviewing our cleaning procedure, we believe the risk of domain specific text cues being present is sufficiently mitigated.
Large Language Models
As a final comparison, we conduct document level bias detection using pre-trained large language models finetuned for chat completion. We experiment with both OpenAI’s GPT Turbo 3.5 (OpenAI 2021) and Meta’s Llama 3 8B Instruct Model (AI@Meta 2024). We pass the following zero-shot prompt template to generate bias predictions: “Please classify the following news article by its political bias. Please only classify the article as {bias labels}, and return no other text. title: {title} article: {article} bias: ”. We do not conduct domain level predictions using the large language models.
Naive Baseline
We use a Zero Rule baseline for a naive model standard, where the test set predictions are all labeled as the majority class in the training set.
4 Results and Discussion
We ran 1284 experiments to train multiple embeddings and supervised learning models over the various configurations and four datasets, to include a topic agnostic dataset that merges the AFG, OATH, and VAX datasets, separate from BASIL. The experiments are trained on 1 node of an NVIDIA DGX A100 system. Each graph autoencoder embedding model takes at most 2 GPU hours to train one model, while each Graph2Vec embedding model takes at most 3.5 CPU minutes. In order to control for different configurations, we formally evaluate the difference between configurations and embeddings using a one-sided Wilcoxan signed rank test on the accuracy and macro F-1 scores. As each pairwise configuration uses the same test sets but are trained independently, the necessary assumptions for a valid hypothesis test are satisfied. This test allows us to determine if there is a paired difference between two methods when all other configurations are held constant.
Through this hypothesis test, We are able to confirm that our graph based embeddings are valuable in comparison to the naive baseline for both metrics, with p-values close to 0. Our experiments have an average .21 increase in macro F-1 scores over the baseline. This is on average a 120% improvement. For the sake of brevity, we display the average and standard deviation of scores for specific configurations in the following tables.
Accuracy Macro F-1
Model
naive baseline 0.522 (.159) 0.246 (.133)
graph2vec_no_features 0.498(.169) 0.341 (.137)
GPT 3.5 Turbo 0.417 (.164) 0.349 (.177)
VGAEautoencode_sent 0.475 (.143) 0.358 (.128)
VGAEautoencode 0.470 (.155) 0.362 (.135)
Llama 3-8B 0.454 (.225) 0.378 (.204)
GAEautoencode_sent 0.524 (.146) 0.391 (.135)
GAEautoencode 0.538 (.146) 0.425 (.138)
doc2vec_dbow_graph2vec_no_features 0.559 (.169) 0.457 (.170)
doc2vec_dm_graph2vec_no_features 0.544 (.178) 0.467 (.184)
doc2vec_dm 0.539 (.170) 0.472 (.170)
doc2vec_dbow 0.574 (.170) 0.486 (.170)
doc2vec_dm_graph2vec_nodewords 0.583 (.164) 0.501 (.167)
graph2vec_spacy 0.589 (.154) 0.501 (.161)
doc2vec_dbow_graph2vec_spacy 0.585 (.163) 0.501 (.165)
graph2vec_nodewords 0.592 (.151) 0.507 (.157)
doc2vec_dm_graph2vec_spacy 0.572 (.157) 0.509 (.160)
doc2vec_dm_graph2vec_nodewords 0.570 (.164) 0.511 (.167)
SBERT 0.607 (.154) 0.542 (.162)
Table 3: Accuracy and Macro F-1 given various embedding models.
Avg (std) of all data configurations
Label Aggregation
We identify that, as expected, there is a large drop in performance as the number of classes increase, with the models trained for fine grained bias classification contributing to the lower overall average scores. The results from models grouped by various label bins are in Table 4. Of note, we find that after regrouping the predictions from models trained using “Full” labels, there is not a statistically significant difference between models trained on the “Full” label set in comparison to “LCR” or “Binary” labels, with p-values of .91 and .36, respectively. This demonstrates that while the “Full” multiclass models have low scores, they are able to extract characteristics of bias in the news’ embedding space just as successfully as the models trained on fewer categories. A one-vs rest model training approach may be used to validate if more granular bias detection is possible with these embeddings.
Accuracy Macro F-1 Baseline Accuracy Baseline Macro F-1
Label Aggregation
Full 0.404 (.102) 0.294 (.111) .380 (.075) .105 (.038)
LCR 0.518 (.102) 0.465 (.102) .470 (.067) .216987 (.036)
Binary 0.722 (.099) 0.608 (.118) .717 (.069) .417 (.024)
Table 4: Accuracy and Macro F-1 given various label aggregation methods.
Avg (std) of all embedding models. The baseline scores are from the naive zero rule baseline
Domain Aggregation
The results for the different domain aggregation comparisons are listed in Table 5. We find that with the BASIL dataset, the embeddings separated more easily into domain level labels (although the article models were still significantly better than the naive baseline across all metrics with a p-value near 0). We also find that the method of aggregating article embeddings makes a significant difference on performance, such that domain models using the Frobenius norm of the difference from its topic perform the worst. The mean of article embeddings across all samples of a domain has the best performance in domain level bias detection. This could both imply that topic information is not significant in domain level bias detection or that a better way of extracting topic level information is warranted.
Accuracy Macro F-1
Domain Aggregation
froNormTopicDiff 0.500 (.159) 0.397 (.153)
no_aggregation 0.565 (.165) 0.453 (.167)
avgTopicDiff 0.556 (.163) 0.483 (.167)
arithmeticMean 0.563 (.168) 0.490 (.175)
BASIL_articles 0.372 (.111) 0.332 (.112)
BASIL_domains 0.588 (.190) 0.544 (.196)
Article Subsets
BASIL 0.480 (.189) 0.438 (.191)
no filter 0.520 (.173) 0.445 (.174)
AFG 0.549 (.157) 0.447 (.143)
VAX 0.510 (.148) 0.451 (.166)
OATH 0.560 (.161) 0.461 (.160)
Table 5: Accuracy and Macro F-1 given additional configurations of domain aggregation and article subsets. Average (std)
Embedding Model
The average mean and standard deviation of scores given different embedding models are in Table 3. Through the paired rank test on the full set of experiments, the GPT, Llama, graph autoencoders, and Graph2Vec with no features (i.e. using the base graph) perform the worst, with no statistically significant performance difference between them. The unsupervised Graph2Vec based embeddings with node attributes are able to capture characteristics of biased text with performance often times similar to, and in the case of “graph2vec_nodewords”, 30% of the time, better than SBERT. However, we find that on average SBERT still outperforms all other embeddings using a p-value significance threshold of .05. Graph2Vec with node attributes (both with node words and word vectors) outperforms all Doc2Vec embedding models with p-values all less than .01. We also observe that when Graph2Vec embeddings with node features are combined with Doc2Vec, a statistically significant improvement (p-values are all less than .01) from the isolated Doc2Vec model occurs. For instance, Graph2Vec with node words concatenated with Doc2Vec with distributed memory (“doc2vec_dm_graph2vec_nodewords”) consistently outperforms embeddings using just Doc2Vec, as observed in their average scores in Table 3. We are not able to identify a statistical difference between embedding performance based on the type of node attribute used in Graph2Vec (node words or Spacy vectors).
Although our graph based embeddings do not consistently outperform SBERT, these results still demonstrate the value of including semantic structure in bias classification. Unlike with SBERT, language models and word embeddings also are unnecessary in this characterization. As demonstrated by the performance of “graph2vec_nodewords”, so long as there is a unique node identifier that can be used to indirectly relate separate graphs, the article graph embeddings are comparably valuable to SBERT. This indirect connection between document networks is necessary and sufficient to have performance that comparable to language model based document embeddings. Without factoring in edge attributes, pre-trained word embeddings, or word frequency outside of missing edges due to non positive PMI, our document networks are able to successfully embed information for bias detection.
On the other hand, even as our computationally expensive graph autoencoder models pass both edge and node attributes in generating document embeddings, they still under perform in comparison to the Graph2Vec unsupervised embeddings. Further experimentation and tuning could be conducted to validate if alternative deep learning graph models, like graph attention or diffusion based approaches (Nikolentzos, Tixier, and Vazirgiannis 2020; Liu et al. 2021), could be repurposed for inductive embedding generation, since graph neural network based approaches are often prone to oversmoothing, or if new graph hashing algorithms that efficiently encode edge attributes could be used in Graph2Vec. Regardless, the fact that a simple approximate graph isomorphism comparison through WL-hashing using minimal additional attributes can generate good bias predictions, validates the value of characterizing semantic structure using document networks.
DocNet Characteristics Case Study
In the following section, we investigate if the models that use Graph2Vec embeddings are extracting distinct graph features that align with bias labels.
To further explore the characteristics of the semantic structure, we obtain descriptive graph metrics for every article and calculate their Pearson correlation with an ordinal encoding of the bias labels as shown in Figures 1 and 2. In both correlation plots, there is not a strong correlation between any of the metrics and the corresponding bias labels. We then train a logistic regression model for each dataset to predict binary bias labels using these metrics. These models have an average .21 lower Macro F-1 on the topic aligned datasets and a .12 lower Macro F-1 on the domain level Basil predictions compared to the corresponding Graph2Vec with node words models. This analysis demonstrates that while the semantic structure of articles holds bias information, there is additional information to be gained by using our proposed graph embeddings rather than curated descriptive graph metrics.
Figure 1: Correlation Plot of Graph Metrics and Bias from Topic Aligned Dataset
Figure 2: Correlation Plot of Graph Metrics and Bias from BASIL (Art is article bias, D is domain bias)
We also plot graphs from the top performing binary graph2vec_node_words embedding models to manually inspect for any consistent characteristics or discrepancies in the data labels. We highlight graphs from the domain label BASIL model, and no domain aggregation VAX, and OATH datasets in Figures 3, 4 and 5, respectively. These graphs are the articles with the highest probability incorrect and correct predictions for both bias labels. The document networks are graphed using a spring force-directed algorithm (Hagberg, Swart, and S Chult 2008).
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 3: Graphs of articles with high probability predictions from BASIL. Predictions from graph2vec_nodewords at the domain level and binary labels (Macro F-1 = .82). Source domains: a) NYTimes, b) Fox, c) Huffington Post d) NYTimes
We observe that in Figure 3, the graphs for the articles that were predicted to be unbiased (a and c) both have visually sparse layouts. Meanwhile, the bias predicted articles (b and d) appear to be denser. Upon returning to the original data, we find that the news article from Figure 3c is actually labeled at the article level to be unbiased. Meanwhile, the original article label for Figure 3d is also labeled to be unbiased. Yet the article corresponding to graph 3d contains phrases like “Donald J. Trump leads a familiar refrain” and “Mr. Trump, who has been criticized for lacing his campaign speeches with profanity”, which some may consider to be inflammatory given word choice like “refrain” and “lacing” (Roappeport 2016). This example highlights the continued difficulty in developing reliable labels for an inherently subjective application. Furthermore, given the BASIL dataset only has three domains, we are not yet able to determine if these visual patterns are a spurious correlation.
To interrogate these patterns further, we look at the graphs of the articles using the article level “graph2vec_node_words” models from the VAX and OATH datasets as shown in Figures 4 and 5. As is apparent in the figures, there is no consistent visual pattern between the articles that are predicted to be unbiased and those that are predicted to be biased in either dataset. This reinforces our finding that a graph embedding approach is needed to capture the complex semantic structure characteristics of biased articles.
While these topic aligned datasets use weak bias labels, there is significantly greater variety in domains in these datasets in comparison to BASIL. This reduces the possibility the bias detection model trains on spurious domain specific characteristics. In validating our weak labels and looking into the article texts from Figures 4 and 5’s graphs, we find a clear usage of informational bias. The article corresponding to 5b ends with the sentence “Those words are directional instructions on Claymore anti-personal mines”(Brigham 2022), perhaps added to invoke a sense of incredulity of the actions of the Oath Keepers. Similarly, the AFN article from 4b opens with the title “Military docs drop vaccine bombshell but ‘fact-check’ has an explanation” (Jordahl and Billy 2022). The usage of “bombshell’ and quotation marks around “fact-check” are strong examples of lexical bias. Both articles align with their domain bias labels. On the other hand, the incorrectly classified OATH MSN article (Figure 5d), which is originally sourced from CNN, contains artifacts like “screaming match”, “The two continued to cut each other off…”, and “chastised”. Although not immediately clear without greater context which partisan lean it has, these informational and lexical rhetorical decisions lead to the assessment that this article is more biased than it is center. If the article had been correctly attributed to CNN in our dataset, it would have been aggregated as biased, rather than the unbiased, specifically left-center, label given to MSN (Rabinowitz 2016). Ultimately, this inspection demonstrates that while there are instances where our labeling assumptions falter, especially in borderline biased articles, our assumption on the generalizability of domain labels to weak article labels holds. Furthermore, our bias detection models may be more astute that these weak labels.
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 4: Graphs of articles with high probability predictions from the VAX dataset. Predictions from graph2vec_nodewords using binary labels (Macro F-1 = .75). Source domains are: a) ABC, b) AFN, c) Toronto Sun d) MSN
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 5: Graphs of articles with high probability predictions from the OATH dataset. Predictions from doc2vec_dbow_graph2vec_nodewords using binary labels (Macro F-1 = .76). Source domains are: a) MSN, b) Raw Story, c) Salon d) MSN
5 Conclusion and Future Work
Through our graph embeddings, we demonstrate that encoding semantic structure in news articles as co-occurrence networks can be used for both domain and article level bias detection. We couple unsupervised embeddings with a logistic regression to learn a separation of the embedding space. We also learn that biased texts have similar context agnostic semantic structures that can be captured by document networks (i.e. DocNet), regardless of the article topic and partisan lean. We show that you can get performance comparable to, and occasionally better than, transformer based language model by the simple inclusion of semantic structure in the form of word-co-occurrence networks. This is demonstrated by the success of the Graph2Vec embeddings in the supervised models, especially as it often has performance comparable to the SBERT embeddings, without the use of word embeddings. DocNet does not require pre-trained language models to generate word embeddings, nor does it require computationally expensive deep learning techniques. The most computationally time consuming portion of DocNet is the initial graph construction of an article. However, this can be parallelized across a corpus of documents for time savings. It also is an inductive approach, allowing for inexpensive and generalizable predictions on new datasets.
Through the varied groupings of bias labels, we identify that there is often insufficient information in the semantic structure of an article for accurate fine-grained bias classification. Future research using one-vs-rest classifiers rather than our generalized multinomial regression may be able to improve upon our results. This is an observation observed by other recent works (Carragher, Williams, and Carley 2024). On the other hand, fuzzier binary classifications can still be sufficiently characterized by semantic structures. This means that right-leaning news articles and left-leaning news articles have significantly more in common than either side of the political spectrum may believe. Although we are not able to pinpoint exactly what this writing style bias is due to the lack of interpretability in our graph embeddings, we can conclude that the characteristics of partisan rhetoric are more than context-dependent.
In future work, it would also be valuable to research embeddings that can fuse word embeddings with semantic structure. While our graph autoencoders perform better than a naive baseline, they are significantly less useful than our Graph2Vec embeddings in bias detection. Context represented in pre-trained language models is valuable to bias detection, as demonstrated in our SBERT model. Future developments in inductive graph embedding approaches that can efficiently and effectively represent dense networks and attributes would likely improve bias detection efforts. In the interim, our DocNet methodology provides a pathway for low-resourced bias detection, where computational resources, rigorously labeled datasets, or language models are not available. These findings may be especially advantageous to low resource languages and communities.
6 Limitations
This project relies on partisan topics and does not look into bias detection of non-partisan capacities. In addition, as the data was collected based on what Tweets on the same topic were shared, there may be a skew in our dataset to highlight media with characteristics that draw the attention of Twitter users. However, given the size of the dataset, the variation in domains present, and the consistency in trends with the manually labeled BASIL dataset, we believe our data to be sufficiently generalizable. Without further data validation and given copyright restrictions, our data should only be used for research purposes. An incorrect bias detection model may lead to misplaced mistrust or over confidence in media. Additionally, this project also only analyzes articles in the English language. While we believe our methodology is language agnostic, this hypothesis has not yet been validated on news articles from other languages.